OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN!!!! Debbie and Trisha are looking at the equation the square root of the quantity of 4 times x minus 6 equals square root of x . Debbie says that the solution is extraneous. Trisha says that the solution is non-extraneous. Is Debbie correct? Is Trisha correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (anonymous):

ok lets write equation on left hand side square root of the quantity of 4 times x minus 6 first write without squareroot 4 times x= 4x and subtract 6 now .....so is 4x-6 now take square root 4x−6−−−−−√

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

they cant be both correct

OpenStudy (anonymous):

now you do the rest

OpenStudy (anonymous):

Okay I'm super confused o_o

OpenStudy (anonymous):

@Flightmaster201

OpenStudy (anonymous):

hey write the equation in numeric form u will understand all

OpenStudy (anonymous):

I'm just really bad at this. It would help if you like walked me through it and tried to teach me.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4 times x means 4 *x which can be wrtten as 4x

OpenStudy (anonymous):

squate root 4 times x minus 6 means sqrt 4x -6

OpenStudy (anonymous):

okayy I get that part.

OpenStudy (anonymous):

squate root 4 times x minus 6 equalls sqrt x means sqrt 4x -6 = sqrt x this is the equation given in the qustion

OpenStudy (anonymous):

get it

OpenStudy (anonymous):

now take squares both sides in order to eliminate the sqrt then the equation will become 4x -6 =x 4x = x + 6 4x - x = 6 3x = 6 divide both sides by 3 x = 2 so which is not extranous roots

OpenStudy (anonymous):

debbie is wrong which means they both r wrong then trisha is right